### v2

# Relevance of long-range screening in Mott transition examined via a hydrogen lattice

Abstract
The Mott transition, a metal-insulator transition due to strong electronic interaction, is observed in many materials without an accompanying change of system symmetry. An important open question in Mott’s proposal is the role of long-range screening, whose drastic change across the quantum phase transition may self-consistently make the transition more abrupt, toward a first-order one. Here we investigate this effect in a model system of hydrogen atoms in a cubic lattice, using charge self-consistent dynamical mean-field theory that incorporates approximately the long-range interaction within the density functional treatment. We found that the system is well within the charge-transfer regime and that the charge-transfer gap intimately related to the Mott transition closes smoothly instead. This indicates that the long-range screening does not play an essential role in this prototypical example. This finding can be understood from the fact that the obtained insulating phase in this model system is driven by strong local interaction, and the transition is associated with the closing of charge-transfer gap. Contrary to Mott’s length scale argument, such energetic competition between kinetic energy and local interaction is thus insensitive to long-range screening.
Mott transition, as an interaction-driven metal-insulator transition (MIT) without assistance of symmetry change, continues to attract a great attention in the past decades [1, 2, 3]. It describes the phase transition between a metal and the Mott insulator widely observed in strongly correlated materials, such as transition-metal oxides [4, 5, 6, 7, 8]. To date, Mott insulators have become one of the most important platforms for the exploration of strongly correlated physics such as unconventional superconductivity [4, 9] and magnetism [10].
The unique characteristic of Mott insulators lies in their unexpected insulating behavior. Traditional insulators have their Bloch orbitals completely filled by electrons such that a finite energy scale is needed for the electrons to be able to propagate without violating the Pauli principle. Mott insulators, on the other hand, do not satisfy this condition and therefore would appear metallic in standard band theories, indicating the need to include additional physics beyond the Pauli principle. To understand such insulating behavior, Mott [11, 1] firstly pointed out the importance of electronic interactions, under which the electrons in the system would reside in a bound state and thus unable to propagate without overcoming the effective binding energy.
Following Mott’s general picture, modern theoretical studies [12, 13] of Mott insulators typically make use of half-filled Hubbard model [14, 15, 16, 17] or its generalized form. For example, transition-metal oxides [6, 7, 8, 4] and organic Mott insulators [18, 19] have relatively strong local electronic repulsion $U$ compared to the kinetic energy. In half-filled systems, such a large $U$ pushes states with doubly occupied orbitals toward high energy and thus allow them to be “integrated out” from the low-energy sector, leaving only effectively singly occupied states without charge freedom (thus an insulator.) From the perspective of Mott’s picture, the repulsion $U$ in such half-filled systems provides exactly the necessary electronic interaction to bind the electron in an effective local orbital.
Still, an important physics question remains open, namely the role of the long-range Coulomb interaction on the nature of the quantum Mott transition. Mott argued [11, 1] that between the metallic and insulating phases the significant change in the screening of long-range interaction could lead to a first-order metal-insulator transition at zero temperature. In contrast, current Hubbard model-based studies (which only incorporate local interaction) found the quantum phase transition to be continuous [20, 12] instead. It is therefore an important and timely task to investigate the role of long-range interaction on the Mott transition, in particular to examine the validity of Mott’s expectation.
Figure 1: Mott’s picture of the key scales of metal-insulator transition: (a) radius $a_{B}$ of the particle-hole bound state, and (b) range $\lambda$ of screened long-range Coulomb attraction $V(r)$ as a function of relative distance $r$. (c) When the bound state density $\rho_{B}$ is confined within the Coulomb attraction, $a_{B}<\lambda$, the system is insulating. (d) Otherwise the system becomes metallic. The finite $\lambda$ at the transition dictates a first-order phase transition in this picture.
Here, to address this essential question, we attempt to incorporate the effect of long-range interaction and investigate the Mott transition of a model system with hydrogen atoms in a cubic lattice. The effect of interaction screening is incorporated approximately via density self-consistency [21, 22] of the density functional theory plus dynamic mean-filed theory (DFT+DMFT) [23, 24, 25, 26, 27, 28, 29, 12, 13, 30]. We found that this model system is in the charge-transfer regime, in which the relevant low-energy charge sector is dominated by the 2ss and singly occupied 1ss orbitals. Correspondingly, the Mott transition is intimately related to the closing of the charge-transfer gap. Importantly, the insulating gap is found to reduce smoothly to zero and the charge carrier density (reflected by $1s$ orbital occupations) grows continuously. Both results suggest a continuous quantum phase transition, instead of the first-order transition expected by Mott’s consideration of long-range screening. Our result can be understood from the fact that this common type of Mott transition in materials is mainly controlled by competition between kinetic and local interaction energies. Contrary to Mott’s length scale argument, such an energetic competition is therefore insensitive to screening of long-range interaction.
[…]
Let’s first review the original consideration of the Mott transition. Mott proposed that a lattice of one-electron atom, such as hydrogen, must be an insulator at large lattice spacing. This is because starting from a ground state with a uniform charge of one electron per atom, the low-energy current response must be gapped associated with the formation of particle-hole bound states with an effective Bohr radius $a_{B}$ [cf. Fig. 1(a)] (assuming a poorly screened $e^{2}/r$ attractive Coulomb interaction over distance $r$.) On the other hand, at small lattice spacing the system is expected to become a good metal with a half-filled lowest energy band. This is because the current response now turns gapless in association with the delocalization of particle-hole pairs, as a result of the long-range screening that reduces the range of the attraction $(e^{2}/r)\text{exp}(-2\pi r/\lambda)$ to a finite length scale $\lambda$ [cf. Fig. 1(b)]. Therefore, in such a length scale consideration, there must exist a transition point for a metal-insulator transition at some intermediate lattice spacing.
As a simple estimation, Mott proposed that the metal-insulator transition would take place at the point when $a_{B}$ and $\lambda$ are equal. This seems quite reasonable since at $a_{B}\ll\lambda$ [cf. Fig. 1(c)] the bound state is well preserved such that the system should be a good insulator, while at $a_{B}\gg\lambda$ [cf. Fig. 1(d)] the interaction is screened too strongly to maintain a bound state such that the system would be a good metal. Based on this rough criteria, Mott estimated the critical density $n_{c}^{1/3}a_{B}\sim 0.25$ [31] via Thomas-Fermi approximation. Though this estimation is very crude, the criteria actually works very well for many materials [2].
The necessity of a finite $\lambda\sim a_{B}$ in Mott’s criterion of metal-insulator transition suggests strongly that the transition be first-order, since a finite $\lambda$ corresponds to a finite itinerant carrier density. As another way to visualize this, near the critical density, a slight increase of itinerant electrons would lead to an enhanced screening (or reduced $\lambda$), which in turn further increases the carrier density. One would therefore expect that such a non-linear feedback can lead to a discontinuous transition at zero temperature. Unfortunately due to the complexity of quantum many-body problems, theoretical studies to date have not been able to incorporate the screening of the length scale of long-range interactions, nor has the corresponding first-order quantum phase transition been obtained [12, 13].
To include this important screening effect of long-range Coulomb interaction beyond the Thomas-Fermi approximation, we use the density self-consistency scheme [21, 22] of density functional theory (DFT) plus dynamical mean-field theory (DMFT) [32, 33, 34, 35, 7, 36], with previously established double counting formulation [22, 30]. In the process of self-consistent iteration, the change of electron density and corresponding screening effect will feedback to the next iteration and thus allow the possibility of the first-order transition. The strong intra-atomic repulsion $U$ ($\sim 1$Ry [30, 37]) is included via DMFT [25, 26, 27, 28, 29, 12, 13] which is so far the most widely applied approximation to achieve the Mott insulating phase. This scheme is implemented by DFT+embedded DMFT functional [21, 22].
[…]
It is obvious that this insulating gap is not the standard Mott gap in a Hubbard model, since its size [38] is only around 22~eV in this case, much smaller than the local repulsion $U\sim 1$~Ry. Indeed, in both DFT [panel (a)] and DFT+DMFT [panel (b)] results, the orbital content right above the gap is predominately of 2ss and 2pp instead of 1ss. Therefore, within the parameter range near the Mott transition, this model system is well within the charge-transfer (CT) insulator regime, in which the lowest-energy charge fluctuation is between 1ss and 2ss/2pp orbitals of neighboring atoms.
Notice that the charge-transfer gap $E_{CT}$ is strongly reduced from 44~eV in the DFT result [panel(a)] to only $\sim 2$~eV in the DFT+DMFT result [panel(b)]. This can be understood from the additional energy lowering from the emergent antiferromagnetic spin exchange between the itinerant carrier in 2ss/2pp and the local immobile 1ss spin, similar to the one that drives the formation of Kondo singlet [39, 40].
Now approaching the metal-insulating transition via reducing the lattice constant, as shown by the black line in Fig. 2(c), the CT gap of the insulating solution reduces and eventually vanishes around $a=a_{c}\sim 2.8$$\text{\AA }$. Naturally, after that the insulator solution can no longer be found and the system is metallic. Incidentally, for $a>a_{c}$ metallic solution is unstable as well. This indicates that the phase transition is intimately related to the gap closing. (Note that in an extremely narrow range of $\Delta a\sim 0.003$$\text{\AA }$ around $a_{c}$, both metallic and insulating solutions can be found [30]. Considerations on this nearly negligible and somewhat artificial feature will be discussed later.)
Most importantly, our results display a very important deviation from Mott’s expectation on the phase transition. Figure 2(c) shows that even with the long-range Coulomb interaction and its screening incorporated via the DFT treatment, as the lattice spacing decreases, $\tilde{E}_{CT}$ only smoothly reduces to zero. Furthermore, as the lattice spacing decreases further, Fig. 2(d) shows that the occupation of the $1s$ orbital (within a fixed atomic sphere [41] of radius 1.8 a.u.) decreases smoothly as well, indicating a gradual development of hole carrier density. No abrupt enhancement of gap closing or the hole carrier density is found at $a_{c}$ that corroborates Mott’s expectation of a first-order phase transition. Clearly, our results suggest a continuous metal-insulator transition. In other words, within the approximation of the state-of-art DFT+DMFT approach (omission of non-local correlation), the proposed drastic effect of screening of long-range interaction does not appear to be essential to the Mott transition in this model system.
This lack of relevance of long-range screening can be understood from the following energetic considerations of this model system. In this charge transfer insulator, mobile carriers can only occur in the ground state when the fully renormalized bandwidths grow to overcome the orbital energy and eliminate the indirect normalized charge-transfer gap minimum, $\tilde{E}_{CT}$, [between momentum ($\pi$$\pi$$\pi$) and ($\pi$,0,0)]. However, this system has a rather large orbital energy difference ($\sim 10$ eV) between the dressed 2s and 1s orbitals and a large kinetic energy represented by the bandwidths of the dressed 1s band ($W_{1s}\sim 5$ eV) and the dressed 2s band ($W_{2s}\sim 20$ eV). Both of these key high-energy factors are insensitive to long-range interaction and thus its screening. Particularly, no dramatic change of long-range screening (of wavevector $q\sim 0$) is expected even right before the indirect gap is closed, since the direct gap remains very large, of 10 eV scale.
In real materials, even beyond this point of gap closing, the mobile electrons in the dressed 2s orbital and holes in the dressed 1s orbital may still be bound into charge-neutral excitons by the long-range interactions. To such an excitonic insulator [42, 43], Mott’s consideration of the insulating phase in Fig. 1(c) may apply, in which the length scale of the long-range interaction is longer than the size of the bound excitons. Nevertheless, given that such excitonic binding should be of the order of 100 meV at best, two orders of magnitude weaker than the above kinetic energies, such an excitonic insulating phase, if present, must only delay the emergence of metallic carriers in a negligibly narrow region. Consequently, only a small number of carriers would be present when entering the metallic phase. Furthermore, this excitonic effect would be greatly suppressed by the dramatic difference in the effective masses of the electrons and holes. Therefore, even if Mott’s proposal of a first-order transition may be realized here, the effect is likely too weak for the accuracy of realistic many-body calculations to date. On the technical side, such excitonic physics that involves non-local particle-hole pairs is surely beyond the capability of the LDA+DMFT framework employed in this study, or even the advocated GW+EDMFT framework [44, 45, 46, 47] as well.
Our conclusion and the analysis above are qualitatively different from those of the previous extended DMFT studies on extended Hubbard model [48, 44, 45, 49, 46, 50, 51, 52, 47, 53, 54, 55], which found a first-order quantum phase transition instead [48]. Note that those results originate from the non-linear feedback in mapping the long-range Coulomb interaction onto retarded (dynamical) screening of local $U$ of the Kondo impurity model. Such mechanism is therefore physically distinct from (and unable to address) Mott’s proposal of screening the length scale of long-range interaction. In contrast, the key factor for the MIT here is the closing of the indirect charge transfer gap, which is insensitive to the long-range screening. Correspondingly, a smooth continuous transition naturally results in our case.
Finally, it is well known [12, 13, 56] that the lack of interatomic correlation in DMFT treatments can lead to an unphysically large degree of degeneracy and in turn an artificially large entropy at finite temperature $T=0^{+}$. This often causes an entropy-driven first-order transition at finite temperature [57, 58]. Our finite-temperature calculation and the effective analytical model above are not immune to this problem [30]. Nonetheless, it is interesting to note that this artifact is much weaker in this model charge-transfer systems than that in the Hubbard model. In fact, in Fig. 2(c), the region hosting first-order transition is narrower than the thickness of the black line. We take it that this distinction reflects the more robust continuous nature of the quantum phase transition in such charge-transfer systems.
Experimentally, our model system of periodic hydrogen lattice could be tested by a new generation of capabilities that engineers hydrogen-like lattice of any desired geometry through STM-implanting donors on surfaces of semiconductors [59]. These experimental advances allow tuning the system across Mott transition without randomness and are thus ideal to experimentally clarify the role of long-range interaction.
In summary, we investigate the important open question concerning the role of long-range screening on the quantum Mott transition using a model system of hydrogen lattice. We employ the charge self-consistent scheme of the dynamical mean-field theory that incorporates approximately the long-range interaction and its screening within the density functional treatment. The system is found to be well within the charge-transfer regime and the charge-transfer gap closes smoothly across the quantum phase transition. Together with a gradually increasing hole carrier density in the metallic phase, this indicates a continuous quantum phase transition distinct from Mott’s proposal. This deviation from Mott’s long-range screening-based first-order picture can be understood from 1) the obtained insulating phase in this typical case is driven by short-range interaction, and 2) the metal-insulator transition involves the closing of the screening-insensitive charge-transfer energy, instead of length-scales switching proposed by Mott. Quantum metal-insulator transitions in such charge-transfer materials are thus insensitive to the long-range interaction. Our study leaves open a finite possibility of weak excitonic insulating phase in a very narrow region, in which Mott’s picture might still apply. Further resolution of this possibility would require a significant advance in current computational capability.
Acknowledgements.
We thank Yuting Tan and Xinyi Li for their helpful discussions. This work is supported by National Natural Science Foundation of China (NSFC) under Grants No. 12274287 & 12042507, and Innovation Program for Quantum Science and Technology No. 2021ZD0301900. S.S. acknowledges support from the Science and Engineering Research Board, India (SRG/2022/000495), (MTR/2022/000638), and IIT(ISM) Dhanbad [FRS(175)/2022- 2023/PHYSICS]. Work in Florida (V.D.) was supported by the NSF Grant No. DMR-2409911, and the National High Magnetic Field Laboratory through the NSF Cooperative Agreement No. DMR-2128556 and the State of Florida.